ThatCreepyBoy:

Solve the triangle shown to the right. (I just need my answer checked, please don't respond until I'm done)

ThatCreepyBoy:

1 attachment
ihy:

1. Find Angle Beta (β) The sum of all angles in a triangle is always 180 degrees. We know two angles: 14.6° and 21.4°. So, β = 180° - 14.6° - 21.4° = 144° Now the next step is to find Side 'a' (Using the Law of Sines) The Law of Sines states: a/sin(A) = b/sin(B) = c/sin(C) And, we can use the known values: b = 18.2, angle B = 144°, and angle A = 21.4° So, a/sin(21.4°) = 18.2/sin(144°) Then, we're gonna rearrange and solve for a: a = (18.2 * sin(21.4°)) / sin(144°) a ≈ 10.98 3. Find Side 'c' (Using the Law of Sines) We can use the same Law of Sines like how we did before We know b = 18.2, angle B = 144°, and angle C = 14.6° So, c/sin(14.6°) = 18.2/sin(144°) Rearrange and solve for c: c = (18.2 * sin(14.6°)) / sin(144°) c ≈ 9.42 In summary: β (Angle Beta): 144° a (Side opposite 21.4°): 10.98 c (Side opposite 14.6°): 9.42

ThatCreepyBoy:

I didn't want you to do it yet, I need to solve this on my own so I can see if I understand how to do it but thank you

ThatCreepyBoy:

Also A and C are incorrect

ihy:

@thatcreepyboy wrote:
Also A and C are incorrect
Hmm..I'll try to figure it out later on

ThatCreepyBoy:

I already got the answer though, I'll show it

ThatCreepyBoy:

1 attachment
ThatCreepyBoy:

1 attachment
ihy:

@thatcreepyboy wrote:
I already got the answer though, I'll show it
Alright, nice job man

ThatCreepyBoy:

1 attachment
ThatCreepyBoy:

thank you

ihy:

@thatcreepyboy wrote:
thank you
Yep, no problem dude